OpenStudy (anonymous):

Titus deposited $1,508.49 in a savings account that earns 4.2% simple interest. What will Titus’s account balance be in 8 months?

OpenStudy (anonymous):

I'm not sure, but. percent for 8 month will be 4.2*(8/12.0) = 2.8, so balance will be $1508.49*1.028 = $1550.72772

OpenStudy (anonymous):

ishalyapin, why do u need to divide by 12? (" 8 month will be 4.2*(8/12.0) = 2.8, so balance ")

OpenStudy (anonymous):

dasDaName, I guess interest rate is 4.2% per year. But If it is per month, you dont' need to divide.

OpenStudy (anonymous):

oh, ok.just tryin to firgure things out